OpenStudy (anonymous):

An airplane reaches its take off velocity of 60 m/s in 30 s starting from rest. Find the time it spends in going from 40 m/s to 60 m/s.

OpenStudy (anonymous):

V= u+at 60=0 + 30a a=2 V=u +at 60=40+ 2t t=10